# RAD: Region-Aware Diffusion Models for Image Inpainting

### v2

@@ 2 Related Works
DDPMs [30, 12] have introduced a novel approach in image generation by employing an iterative denoising process that progressively refines random noise into high-quality images. Unfortunately, DDPMs showed lower image fidelity compared to GANs. To enhance fidelity, various studies [23, 36, 25] have focused on generating high-quality images using diverse datasets such as ImageNet [7], FFHQ [14], and LSUN [41]. Dhariwal and Nichol [23] showed that achieving high log-likelihood on datasets with high diversity, like ImageNet [7], is possible through a hybrid objective. This hybrid objective facilitates learning the variances of the reverse Gaussian transitions that were fixed in DDPM [12]. Meanwhile, Dhariwal et al. [8] introduced ADMs that use auxiliary classifiers to classify the noisy images generated during the reverse process. This simple class-conditioned generation method, known as classifier guidance, mainly enhances the fidelity of generated images by applying strong class conditioning.
@@
Other approaches leveraged pretrained diffusion models for image inpainting/editing by manipulating their generation processes. Some methods [21, 4] manipulated the reverse SDE procedure of a pretrained score-based model [32]. On the other hand, several works [20, 15, 34, 2] utilize ADM [8]. Notably, Lugmayr et al. [20] proposed RePaint, a method utilizing resampling steps to harmonize mask and non-mask regions. Other studies [10, 6, 1, 5] used the stable diffusion model [27]. Couairon et al. [6] proposed DiffEdit, which employs DDIM inversion for image editing by generating masks based on text prompts to preserve backgrounds.
@@ 3.2 Generating Noise Schedules
To mimic the noise-filling process of DDPM in each phase, we use the following strategy: Let $T_{1}$ and $T_{2}$ be the numbers of timesteps for Phases 1 and 2, respectively, i.e., $T_{1}+T_{2}=T$. During each phase, each pixel in or outside the mask is filled in by a scalar noise with variance $\beta_{t}$ ($1\leq t\leq T_{1}$) or $\beta_{t}^{\prime}$ ($1\leq t\leq T_{2}$), respectively, with $10<\beta_{t},\beta_{t}^{\prime}<$ 1. In practice, we use simple linear schedules for $\beta_{t}$ and $\beta_{t}^{\prime}$ as in DDPM. A caveat here is that all pixels must have the same accumulated noise levels after finishing the two phases, which can be satisfied easily by normalizing $\beta_{t}$ and $\beta_{t}^{\prime}$ as explained in the supplementary material. Gathering the above pixel-wise noise schedules, we can form $b_{t}$.
@@ 4.1 Benchmark Datasets
Method FFHQ LSUN Bedroom
Box Extreme Wide Box Extreme Wide
FID $\downarrow$ LPIPS $\downarrow$ FID $\downarrow$ LPIPS $\downarrow$ FID $\downarrow$ LPIPS $\downarrow$ FID $\downarrow$ LPIPS $\downarrow$ FID $\downarrow$ LPIPS $\downarrow$ FID $\downarrow$ LPIPS $\downarrow$
LaMa [33] 27.7 0.086 61.7 0.492 23.2 0.096 - - - - - -
Score-SDE [32] 30.3 0.135 48.6 0.488 29.8 0.132 23.7 0.648 24.1 0.648 23.2 0.644
DDRM [15] 28.4 0.109 48.1 0.532 27.5 0.113 20.5 0.166 33.1 0.450 26.4 0.190
RePAINT [20] 25.7 0.093 35.9 0.398 24.2 0.108 20.5 0.176 23.5 0.461 21.4 0.161
MCG [4] 23.7 0.089 30.6 0.366 22.1 0.099 19.9 0.131 22.0 0.395 20.9 0.108
DDNM [34] 30.4 0.089 87.7 0.353 30.4 0.089 22.7 0.150 53.3 0.431 23.2 0.126
DeqIR [2] 24.2 0.093 64.2 0.368 27.4 0.099 22.2 0.176 43.9 0.461 22.0 0.153
RAD (ours) 22.1 0.074 33.4 0.317 21.5 0.078 19.2 0.131 21.6 0.399 20.8 0.107
@@ 4.2 Implementation Details
For all experiments, we set the rank of LoRA [13] to 16 and trained RAD using the Adam optimizer [16] with a learning rate of $10^{-4}$. We used a batch size of 16 for both FFHQ and ImageNet, while that was eight for LSUN Bedroom. The total diffusion steps was 2000, with the initial 1000 steps dedicated to Phase 1. We used predefined linear noise schedules for $\beta_{t},\beta_{t}^{\prime}\in\left(0.0001,0.02\right)$. During inpainting, we used 100 sampling steps. FFHQ was trained on four TITAN RTX GPUs, LSUN Bedroom on eight NVIDIA 3090 GPUs, and ImageNet on eight NVIDIA RTX 6000 ADA GPUs. Pretrained ADM models [8] were used for all datasets. RAD was trained for 300K on FFHQ and LSUN Bedroom, and 800K on ImageNet.
@@ 4.3 Results
Table 2: Inference time comparison (1×1$\times$1 × NVIDIA TITAN RTX).

### v4

@@ 2 Related Works
DDPMs [31, 13] have introduced a novel approach in image generation by employing an iterative denoising process that progressively refines random noise into high-quality images. Unfortunately, DDPMs showed lower image fidelity compared to GANs. To enhance fidelity, various studies [24, 37, 26] have focused on generating high-quality images using diverse datasets such as ImageNet [7], FFHQ [15], and LSUN [42]. Dhariwal and Nichol [24] showed that achieving high log-likelihood on datasets with high diversity, like ImageNet [7], is possible through a hybrid objective. This hybrid objective facilitates learning the variances of the reverse Gaussian transitions that were fixed in DDPM [13]. Meanwhile, Dhariwal et al. [8] introduced ablated diffusion models (ADMs) that use auxiliary classifiers to classify the noisy images generated during the reverse process. This simple class-conditioned generation method, known as classifier guidance, mainly enhances the fidelity of generated images by applying strong class conditioning.
@@
Other approaches leveraged pretrained diffusion models for image inpainting/editing by manipulating their generation processes. Some methods [22, 4] manipulated the reverse SDE procedure of a pretrained score-based model [33]. On the other hand, several works [21, 16, 35, 2, 10] utilize ADM [8]. Notably, Lugmayr et al. [21] proposed RePaint, a method utilizing resampling steps to harmonize mask and non-mask regions. Other studies [11, 6, 1, 5] used the stable diffusion model [28]. Couairon et al. [6] proposed DiffEdit, which employs DDIM inversion for image editing by generating masks based on text prompts to preserve backgrounds.
@@ 3.2 Generating Noise Schedules
To mimic the noise-filling process of DDPM in each phase, we use the following strategy: Let $T_{1}$ and $T_{2}$ be the numbers of timesteps for Phases 1 and 2, respectively, i.e., $T_{1}+T_{2}=T$. During each phase, each pixel in or outside the mask is filled in by a scalar noise with variance $\beta_{t}$ ($1\leq t\leq T_{1}$) or $\beta_{t}^{\prime}$ ($1\leq t\leq T_{2}$), respectively, with $0<\beta_{t},\beta_{t}^{\prime}<1$. In practice, we use simple linear schedules for $\beta_{t}$ and $\beta_{t}^{\prime}$ as in DDPM. A caveat here is that all pixels must have the same accumulated noise levels after finishing the two phases, which can be satisfied easily by normalizing $\beta_{t}$ and $\beta_{t}^{\prime}$ as explained in the supplementary material. Gathering the above pixel-wise noise schedules, we can form $b_{t}$.
@@ 4.1 Benchmark Datasets
Method FFHQ LSUN Bedroom ImageNet
Box Extreme Wide Box Extreme Wide Box Extreme Wide
FID $\downarrow$ LPIPS $\downarrow$ FID $\downarrow$ LPIPS $\downarrow$ FID $\downarrow$ LPIPS $\downarrow$ FID $\downarrow$ LPIPS $\downarrow$ FID $\downarrow$ LPIPS $\downarrow$ FID $\downarrow$ LPIPS $\downarrow$ FID $\downarrow$ LPIPS $\downarrow$ FID $\downarrow$ LPIPS $\downarrow$ FID $\downarrow$ LPIPS $\downarrow$
LaMa [34] 27.7 0.086 61.7 0.492 23.2 0.096 - - - - - - - - - - - -
Score-SDE [33] 30.3 0.135 48.6 0.488 29.8 0.132 23.7 0.648 24.1 0.648 23.2 0.644 57.2 0.200 86.6 0.495 62.0 0.183
DDRM [16] 28.4 0.109 48.1 0.532 27.5 0.113 20.5 0.166 33.1 0.450 26.4 0.190 74.3 0.224 106.9 0.492 75.1 0.211
RePAINT [21] 25.7 0.093 35.9 0.398 24.2 0.108 20.5 0.176 23.5 0.461 21.4 0.161 54.0 0.177 84.0 0.479 59.0 0.166
MCG [4] 23.7 0.089 30.6 0.366 22.1 0.099 19.9 0.131 22.0 0.395 20.9 0.108 48.1 0.132 58.4 0.448 56.9 0.124
DDNM [35] 30.4 0.089 87.7 0.353 30.4 0.089 22.7 0.150 53.3 0.431 23.2 0.126 63.8 0.187 80.6 0.476 64.5 0.167
DeqIR [2] 24.2 0.093 64.2 0.368 27.4 0.099 22.2 0.176 43.9 0.461 22.0 0.153 66.5 0.195 99.5 0.505 68.4 0.182
RAD (ours) 22.1 0.074 33.4 0.317 21.5 0.078 19.2 0.131 21.6 0.399 20.8 0.107 47.0 0.118 57.8 0.374 56.7 0.104
@@ 4.2 Implementation Details
For all experiments, we set the rank of LoRA [14] to 16 and trained RAD using the Adam optimizer [18] with a learning rate of $10^{-4}$. We used a batch size of 16 for both FFHQ and ImageNet, while that was eight for LSUN Bedroom. The total diffusion steps was 2000, with the initial 1000 steps dedicated to Phase 1. We used predefined linear noise schedules for $\beta_{t},\beta_{t}^{\prime}\in\left(0.0001,0.02\right)$. During inpainting, we used 100 sampling steps. FFHQ was trained on four TITAN RTX GPUs, LSUN Bedroom on eight NVIDIA 3090 GPUs, and ImageNet on eight NVIDIA RTX 6000 ADA GPUs. Pretrained ADM models [8] were used for all datasets. RAD was trained for 300K on FFHQ and LSUN Bedroom, and 700K on ImageNet. ImageNet was trained using an exponential moving average (EMA) with the decay factor set to 0.999.
@@ 4.3 Results
Table 2: Inference time comparison ($1\times$ NVIDIA TITAN RTX).
@@ 6 Conclusion
Acknowledgments
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.RS-2022-00155885, Artificial Intelligence Convergence Innovation Human Resources Development (Hanyang University ERICA)) and partly by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (RS-2023-NR076576).